# Understanding intra-node communication in HPC systems and Datacenters

### v1

@@ 5 Conclusions
This paper has analyzed the impact of the interference between intra- and inter-node communication operations generated by emerging applications (e.g., generative AI or LLMs) in Supercomputers and Datacenters. To help with this research, we have developed a detailed packet-level simulation model validated against a real cluster infrastructure using standard micro-benchmarks, demonstrating its accuracy and reliability. Additionally, we conducted scale-out simulation experiments to model large heterogeneous computing and storage clusters, scaling up to 1024 end-node devices (e.g., accelerators). We have also modeled realistic communication patterns based on LLM programming model using different types of parallelism (data, pipeline and tensor parallelism), which have been used to feed our simulation tool. Our analysis has identified a critical bottleneck at the interface between the intra-node network and the NIC link, which significantly impacts the performance of both intra- and inter-node networks and reduces overall system performance. We have also observed that increasing or decreasing the number of nodes in a supercomputer does not fundamentally alter this bottleneck’s impact, as the intra-node network’s limitations remain the dominant factor. Our simulation and the findings related to intra- and inter-node communication dynamics provide valuable insights to improve the design of large-scale heterogeneous computing environments.

### v2

@@ 5 Conclusions
This paper has analyzed the impact of the interference between intra- and inter-node communication operations generated by emerging applications (e.g., generative AI or LLMs) in Supercomputers and Data Centers. To help with this research, we have developed a detailed packet-level intra- and inter-node network simulation model validated against a real cluster infrastructure using standard micro-benchmarks. We have thoroughly analyzed the latency and overhead effects of inter-to-intra-node packet conversion when different payload sizes are used. Additionally, we conducted scale-out simulation experiments modeling intra- and inter-node networks up to $4096$ accelerators. We have also modeled realistic communication patterns based on the LLMs programming model, using different types of parallelism (data, pipeline, and tensor parallelism), which have been used to feed our simulation tool. Our analysis has identified a critical bottleneck at the interface between the intra-node network and the NIC link, which significantly impacts the performance of both intra- and inter-node networks and reduces overall system performance. We have also observed that increasing or decreasing the number of server nodes or increasing the speed of intra-node links does not alter this bottleneck’s impact, as the intra-node network’s limitations remain the dominant factor. Our simulation and the findings related to intra- and inter-node communication dynamics provide valuable insights to improve the design of large-scale intra- and inter-node interconnection networks.
[…]
Acknowledgements
This work is supported by the Spanish Ministry of Science and Universities MCIN/AEI/10.13039/501100011033, the European Union (NextGenertionEU/PRTR) under project TED2021-130233B-C31, the Junta de Comunidades de Castilla-La Mancha and FEDER funds under the project SBPLY/21/180501/000248, the PERTE-Chip grants (UCLM Chair, TSI-069100-2023-0014) funded by the Spanish Ministry of Digital Transformation and Public Service, and the Universidad de Castilla-La Mancha under project 2023-GRIN-34056.